# Contextual modulation of language comprehension in a dynamic neural model of lexical meaning

### v1

@@ 1.4 Dynamic Field Theory (DFT)
In this subsection, we give a brief overview of DFT; a more detailed description of the relevant mathematics is given in Section 2. In DFT, features relevant for cognition are modeled as continuous parameters governed by the activity of populations of neurons. The activity of a neural population over time is described using a differential equation defining a dynamic neural field (DNF: Amari, 1977). DNFs are characterized by point attractor dynamics. This means that, at any given time, the activation pattern in a DNF is attracted to a particular state, i.e., the point attractor state. The location of the point attractor can change over time according to a variety of factors, e.g., inputs to the DNF from sensory surfaces or other DNFs. Usually, when a DNF is not receiving any input, the point attractor corresponds to an inactive or resting state. When a DNF begins to be influenced by input, the point attractor might shift to an active state. In particular, the dynamics of lateral interaction within DNFs allow the formation of “peaks” of activation. Depending on the cognitive dimension being represented by the DNF, an activation peak might correspond to a movement goal, a percept, or another kind of cognitive event. Since change within DNFs is characterized by point attractor dynamics, but the location of the point attractor is itself ever-changing, DNFs can be described as metastable (e.g., Kelso, 2012). Metastability, invoking the interplay between stability and flexibility, is an important property of cognition (Kelso, 1995). Stability offers resistance to the ubiquitous influence of noise, and flexibility allows rapid change under changing cognitive and environmental conditions. In the context of lexical meaning, metastability offers a way to reconcile intuitions of discreteness with evidence for underlying continuity.
@@
DFT originally developed in the context of motor control research, especially in the domains of eye movements (Kopecz and Schöner, 1995) and arm movements (Erlhagen and Schöner, 2002). It has been increasingly applied in other cognitive domains, including, recently, speech and language. DFT models of speech and language have, so far, focused on the neurocognitive representation of phonetic dimensions, e.g., voice onset time (VOT), and the location and degree of constrictions formed by the tongue. Thus, activation peaks correspond to articulatory movement goals. These models have offered novel explanations for a variety of empirical phenomena: effects of auditory perception on verbal response times (Roon and Gafos, 2016), effects of lexical competitors on speech articulation in errors (Stern et al., 2022) and non-errors (Stern and Shaw, 2023a, b), long-term phonological change (Gafos and Kirov, 2009; Shaw and Tang, 2023), and individual differences in phonological representations (Harper, 2021).
@@ 1.5 This paper
In this study, we extend neural field dynamics to dimensions of semantic interpretation. While the relevant cognitive dimensions are specific to the meaning system and therefore differ from those of previous DFT models of speech, the basic dynamics of neural activity are the same. The model we propose implements a mapping between the lexical item have and the continuous semantic space schematized in Figures 1 and 2. Instead of articulatory movement goals, peaks of neural activation correspond to semantic interpretations. One contribution of the paper is to offer a neurocognitive process explanation of the results described above: contextual modulation of the timecourse and outcome of lexical interpretation, and individual variation in the magnitude of this modulation. More generally, we offer an explanation for the intuition that meaning is experienced as discrete (e.g., adjacency or possession), despite evidence for a continuous substrate. Such apparent discreteness, we propose, arises from metastability in a continuous space defined by neurocognitive variables. In addition to capturing existing results, simulations from the model also generate a novel empirical prediction, which we test with an experiment combining self-paced reading and acceptability judgments. We intend our model to be general, using the English lexical item have as a test case or proof of concept. We propose that, while individual lexical items vary in the relevant semantic dimensions, as well as in the details of their coupling to those dimensions, the basic architecture and mechanisms of our model do not vary.
The remainder of the paper is structured as follows. Section 2 describes the structure of the model and the results of the simulations. Section 3 describes the design of the experiment, and Section 4 reports the experimental results. Section 5 discusses theoretical implications of the study and new predictions that remain to be tested.
@@ 2.1 Model structure
The rate of change of activation is negatively related to current activation , defining a dynamical system with a point attractor at . represents external input to the node, e.g. from perception or intention, and represents normally distributed noise weighted by a parameter . Thus, when there is external input to the node, the node’s activation is attracted to . When there is no input, activation is attracted to 0. is a time constant, with higher values corresponding to slower rates of evolution. In the simulations reported below, is set to , and is set to . We set the magnitude of external input depending on the condition being simulated.
@@
Activation is defined for each neuron representing the relevant semantic dimension at each moment in time . The range of is set to 100, such that each neuron in the field is maximally sensitive (“tuned”) to a particular percentage of the maximum conceivable value of that dimension. For instance, when the neuron in the conn DNF is active, this corresponds to an interpretation of “ of maximum conceivable connectedness”. Activation in the DNF has a point attractor at . The resting level is assumed to be below zero for all neurons, by convention at . We set the time constant for the DNFs to . Each field input , , and is represented as a separate Gaussian distribution of the form
@@
The have node is coupled to the conn DNF with a wide distribution ( = 40) positioned at the center of the field ( = 50), as seen in Figure 3. When the have node becomes active, it sends input to the entire conn field, consistent with an analysis of polysemy for the lexical semantics of have. However, the center of the conn field is favored, consistent with the fact that, all else equal, have tends to evoke an interpretation of alienable possession (intermediate connectedness).
represents input from each DNF to the other. Via this mechanism, patterns of activation in one DNF evoke corresponding patterns of activation in the other DNF. This mechanism implements the “inverted U” pattern described in Section 1 (Figure 2). In particular, as seen in Figure 4, activation consistent with high control asymmetry evokes activation consistent with intermediate connectedness (and vice versa; note the double-sided arrows), and activation consistent with low control asymmetry evokes activation consistent with both low connectedness and high connectedness (and vice versa).
@@
The term on the right, , defines the basic magnitude of . The magnitude of input from each neuron in the field which is sending input (of size ) to the field receiving input is determined by its activation above resting level , weighted by a parameter , set to 0.3 in our simulations. This magnitude is further weighted by , a nonlinear (quartic) function of the distance from to the center of the sending distribution , divided by the width of the sending distribution . Neurons within one of contribute substantially to , while neurons exceeding one from contribute exponentially less. This value is summed for all neurons within the sending field, and then weighted by the term on the left, . When the maximum activation within the sending distribution (ranging from to ) is much greater than the maximum activation in the receiving distribution, this term approaches 1, so . However, as the maximum activation in the receiving distribution approaches (or exceeds) the maximum activation in the sending distribution, the weighting term approaches 0, reducing or eliminating . In this way, the sending field cannot increase activation in the receiving field beyond its own maximum activation, preventing an infinite positive feedback loop. The centers and widths of each distribution are given in Table 1.
@@ 2.2 Model simulations
In this section, we use the model to simulate the interpretation of sentences conveying adjacency meanings of have in two contexts: (1) following an adjacency interpretation of have, and (2) following a possession interpretation of have. Each simulation consists of three phases, summarized in Table 2. In phase 1, the model receives external inputs (all with amplitude ) which drive interpretation of the context sentence. In the adjacency context, these inputs correspond to low CA and low conn; the have node also receives input. In the possession context, external inputs also excite the have node and the low side of the CA field, but the conn field receives input corresponding to high connectedness, i.e., inalienable possession. Phase 1 is the only phase which differs between conditions. In phase 2, all inputs are removed. This corresponds to the time between interpreting the context sentence and interpreting the target sentence. In phase 3, the model receives external inputs corresponding to the adjacency have target sentence. These inputs are identical in both conditions: have and low CA. Phases 1 and 3 each run for 90 timesteps, which we found to be enough time for an activation peak to stabilize in both DNFs. Phase 2 runs for 20 timesteps, which we found to be enough time for both activation peaks to fall below the interaction threshold (), but not enough time for the fields to return fully to the resting level. Each simulation runs for a total of 200 timesteps.
@@
| adjacency context | low CA low conn have | - | low CA have |
| possession context | low CA high conn have | - | low CA have |
@@
Examples of simulated interpretation in each context are displayed in Figure 7. In phase 1 (up to timestep 90) the have node becomes active, and begins to send input to the conn field. The CA field quickly forms a stabilized peak corresponding to an interpretation of low control asymmetry in both conditions. Due to field coupling, this CA peak sends input to the conn field on both the low and high ends. In the adjacency context, an activation peak corresponding to low conn stabilizes and suppresses the rest of the field via lateral inhibition. In the possession context, the winning activation peak is on the high side of the field. During phase 2 (from timesteps 91–110), all activation decreases towards resting level. Activation of the have node returns almost back to its resting level. In the fields, activation falls below the interaction threshold (), but does not return fully to the resting level. At timestep 111 (the beginning of phase 3), field activation is still highly non-uniform, reflecting residual activation from phase 1. In the CA field, another peak on the low side of the field forms in both conditions. In the conn field, the location of the peak differs by condition, despite the fact that the phase 3 inputs are identical between conditions. In the adjacency context, phase 3 (target) processing stabilizes on another interpretation of low conn; in the possession context, the stabilized peak corresponds to an interpretation of high conn. In this way, the evolution of activation is shaped by the preceding state of the system.
@@
The examples in Figure 7 were selected in order to demonstrate the effect of preceding context (phase 1 evolution) on target sentence interpretation (phase 3 evolution). However, not every simulated run of the model exhibits the same effect. The presence of noise in the model introduces a stochastic influence on the location of field stabilization in each simulation. In order to examine the robustness of the contextual modulation effect, we simulated 1000 instances of interpretation in each of the two conditions. As seen in Figure 8, there is a bimodal distribution of interpretations in both contexts. That is, for each context, an interpretation of either adjacency or possession for the target sentence was possible. However, the likelihood of each interpretation was influenced by context. In the adjacency context, low conn (adjacency) interpretations were more likely; in the possession context, high conn (possession) interpretations were more likely. Thus, while context does not completely determine the course of field evolution, it exerts a strong enough influence to be observable over many simulations.
How do we relate these simulated results to the observed acceptability results described in Section 1? Field stabilization at the low end of the conn field corresponds to an adjacency interpretation, which is consistent with the two inanimate arguments of have in the target sentence. Thus, a low conn peak should correlate with a higher acceptability rating. On the other hand, a high conn peak, corresponding to a possession interpretation, is inconsistent with the two nouns in the target sentence, which cannot easily be construed to be in a possession relationship. Thus, a high conn peak should correlate with a lower acceptability rating. In order to transform the field stabilization results into simulated acceptability results, we therefore reverse the value of the location of field stabilization using the following equation
where ranges from to . As seen in Figure 9, simulated acceptability was higher in the adjacency context relative to the possession context, consistent with existing results.

### v2

@@ 1.5 Dynamic Field Theory (DFT)
In this subsection, we give a brief overview of DFT; a more detailed description of the relevant mathematics is given in Section 2. In DFT, features relevant for cognition are modeled as continuous parameters governed by the activity of populations of neurons. The activity of a neural population over time is described using a differential equation defining a dynamic neural field (DNF: Amari, 1977). DNFs are characterized by point attractor dynamics. This means that, at any given time, the activation pattern in a DNF is attracted to a particular state, i.e., the point attractor state. The location of the point attractor can change over time according to a variety of factors, e.g., inputs to the DNF from sensory surfaces or other DNFs. Since change within DNFs is characterized by point attractor dynamics, but the location of the point attractor changes over time, DNFs are characterized by an interplay between stability and flexibility. Stability offers resistance to the ubiquitous influence of noise, and flexibility allows rapid change under changing cognitive and environmental conditions (e.g., Kelso, 1995). Usually, when a DNF is not receiving any input, the point attractor corresponds to an inactive or resting state. When a DNF begins to be influenced by input, the point attractor might shift to an active state. In particular, the dynamics of lateral interaction within DNFs allow the formation of “peaks” of activation. Depending on the cognitive dimension being represented by the DNF, an activation peak might correspond to a movement goal, a percept, or another kind of cognitive event. Discontinuous shifts from inactive states to active states (and vice versa) exemplify the nonlinear dynamics of DNFs, whereby continuous change in one variable (e.g., input to a DNF) can result in sudden, discontinuous change in another variable (e.g., DNF activation). In the context of lexical meaning, nonlinear dynamics offer a way to reconcile intuitions of discreteness with evidence for underlying continuity.
1.5.1 DFT and language
@@
DFT originally developed in the context of motor control research, especially in the domains of eye movements (Kopecz and Schöner, 1995) and arm movements (Erlhagen and Schöner, 2002). It has been increasingly applied in other cognitive domains, including, recently, speech and language. A number of DFT models of speech and language have focused on the neurocognitive representation of phonetic dimensions, e.g., voice onset time (VOT), and the location and degree of constrictions formed by the tongue. In these models, activation peaks correspond to articulatory movement goals. These models have offered novel explanations for a variety of empirical phenomena: effects of auditory perception on verbal response times (Roon and Gafos, 2016), effects of lexical competitors on speech articulation in errors (Stern et al., 2022) and non-errors (Stern and Shaw, 2023a, b), long-term phonological change (Gafos and Kirov, 2009; Kirkham and Strycharczuk, 2024; Shaw and Tang, 2023), and individual differences in phonological representations (Harper, 2021). Other DFT models of language have focused on the processing of words and phrases describing physical properties of objects like color, size, and spatial position (Bhat et al., 2022; Kati et al., 2024; Richter et al., 2021; Sabinasz et al., 2023). DFT is particularly useful for unifying discrete and continuous linguistic representations, and synthesizing empirical results and theoretical insights from different subfields in the study of human language (Stern, 2025).
@@ 1.6 This paper
In this study, we extend neural field dynamics to semantic dimensions underlying lexical polysemy. The model we propose implements a mapping between the lexical item have and the continuous semantic space schematized in Figures 1 and 2. In the model, peaks of neural activation correspond to semantic interpretations. One contribution of the paper is to offer a neurocognitive process-based explanation of the behavioral results described above: contextual modulation of the timecourse and outcome of lexical interpretation, and individual variation in the magnitude of this modulation. Our explanation captures the intuition that meaning is experienced as discrete (e.g., adjacency or possession), despite evidence for a continuous substrate. Such apparent discreteness, we propose, arises from nonlinear dynamics of continuous neurocognitive variables. In addition to capturing existing results, simulations from the model also generate a novel empirical prediction, which we test with an experiment combining self-paced reading and acceptability judgments. We intend our model to be general, using the English lexical item have as a test case. We propose that, while individual lexical items vary in the relevant semantic dimensions, as well as in the details of their coupling to those dimensions, the basic architecture and mechanisms of our model do not vary.
The remainder of the paper is structured as follows. Section 2 describes the structure of the model and the results of the simulations. Section 3 describes the design and results of the behavioral experiment. Section 4 discusses theoretical implications of the study and new predictions that remain to be tested.
@@ 2.1 Model structure
The rate of change of activation is negatively related to current activation , defining a dynamical system with a point attractor at . represents external input to the node, e.g. from perception or intention, and represents normally distributed noise weighted by a parameter . When there is external input to the node, the node’s activation is attracted to . When there is no input, activation is attracted to 0 . is a time constant, with higher values corresponding to slower rates of evolution. In the simulations reported below, is set to , and is set to . We set the magnitude of external input depending on the condition being simulated.
@@
Activation is defined for each neuron representing the relevant semantic dimension at each moment in time . The range of is set to 99, such that each neuron in the field is maximally sensitive (“tuned”) to a particular percentage of the maximum conceivable value of that dimension. For instance, when the neuron in the conn DNF is active, this corresponds to an interpretation of “ of maximum conceivable connectedness”. Activation in the DNF has a point attractor at . The resting level is assumed to be below zero for all neurons, by convention at . We set the time constant for the DNFs to . Each field input , , and is represented as a separate Gaussian distribution of the form
@@
The have node is coupled to the conn DNF with a wide distribution ( = 40) positioned at the center of the field ( = 50), as seen in Figure 3. When the have node becomes active, it sends input to the entire conn field, consistent with an analysis of polysemy for the lexical semantics of have. However, the center of the conn field is favored, consistent with the fact that, all else equal, have privileges an interpretation of alienable possession (intermediate connectedness), as described in Section 1.1.
represents input from each DNF to the other. Via this mechanism, patterns of activation in one DNF evoke corresponding patterns of activation in the other DNF. This mechanism implements the downward parabola pattern described in Section 1 (Figure 2). In particular, as seen in Figure 4, activation consistent with high control asymmetry evokes activation consistent with intermediate connectedness (and vice versa; note the double-sided arrows), and activation consistent with low control asymmetry evokes activation consistent with both low connectedness and high connectedness (and vice versa).
@@
The term on the right, , defines the basic magnitude of . The magnitude of input from each neuron in the field which is sending input (of size ) to the field receiving input is determined by its activation above resting level , weighted by a parameter , set to in our simulations. This magnitude is further weighted by , a nonlinear (quartic) function of the distance from to the center of the sending distribution , divided by the width of the sending distribution . Neurons within one of contribute substantially to , while neurons exceeding one from contribute exponentially less. This value is summed for all neurons within the sending field, and then weighted by the term on the left, . When the maximum activation within the sending distribution (ranging from to ) is much greater than the maximum activation in the receiving distribution, this term approaches 1, so . However, as the maximum activation in the receiving distribution approaches (or exceeds) the maximum activation in the sending distribution, the weighting term approaches 0, reducing or eliminating . In this way, the sending field cannot increase activation in the receiving field beyond its own maximum activation, preventing an infinite positive feedback loop. The centers and widths of each distribution are given in Table 1.
@@ 2.2 Model simulations
In this section, we use the model to simulate interpretation of sentences containing have. The main purpose of the simulations is to investigate how context influences interpretation of anomalous sentences like 1.1, analogous to the experiment from Zhang et al. (2018), described in Section 1.3. We present these simulation in Sections 2.2.2 and 2.2.3. First, in Section 2.2.1, we demonstrate how the model interprets a more canonical use of have, as in 2.2.
. The professor has a motorcycle.
2.2.1 Example simulation: alienable possession
To simulate interpretation of the sentence in 2.2, the model receives three external inputs, all with : input to the have node, input corresponding to high CA, and input corresponding to mid conn. The input to the have node corresponds to perception of the lexical item have. The other two inputs correspond to knowledge about the referents of the nouns serving as arguments of have: professor and motorcycle. We simulated model evolution under the influence of these three inputs for 90 timesteps. Figure 7 displays the results.
It can be seen that all three model components stabilize in active states shortly after the onset of the simulation. The have node is attracted to because of the external input to the node. The CA field forms an activation peak at the high end of the field because of external input centered at this location, as well as input from the conn field centered at this location (see Figure 4). The conn field forms an activation peak in the middle of the field because of external input, input from the have node, and input from the CA field. This qualitative activation state will persist until the inputs change or are removed. In this example, the model stabilizes quickly, and predictably, because the various influences on activation—external inputs, node-field coupling, and field-field coupling—reinforce each other.
2.2.2 Example simulations: contextual modulation
Next, we simulate sentences containing have with two unrelated, inanimate arguments (as in sentence 1.1) in two contexts: following an adjacency reading of have, and following a possession reading of have. These simulations are analogous to the experiment in Zhang et al. (2018), described in Section 1.3. Each simulation consists of three phases, summarized in Table 2. In phase 1, the model receives external inputs which drive interpretation of the context sentence. In the adjacency context (corresponding to sentence 1.1), these inputs correspond to low CA and low conn because of the meanings of the nouns and the locative prepositional phrase; the have node also receives input. In the possession context (corresponding to sentence 1.3), external inputs also excite the have node and the low side of the CA field, but the conn field receives input corresponding to high connectedness, i.e., inalienable possession. Again, the latter two inputs come from the meanings of the nouns which are the arguments of have. Phase 1 is the only phase which differs between conditions. In phase 2, all inputs are removed. This corresponds to the time between interpreting the context sentence and interpreting the target sentence. In phase 3, the model receives external inputs corresponding to the target sentence. These inputs are identical in both conditions: have (), low CA (), and low conn (). The amplitude of the input to the low end of the conn field is much weaker in the target sentence relative to the adjacency context sentence, because the target sentence lacks a locative prepositional phrase. Nonetheless, the two unrelated, inanimate nouns in the target sentence are assumed to induce a weak bias for a low connectedness interpretation. Phases 1 and 3 each run for 90 timesteps, which was found to be enough time for an activation peak to stabilize in both DNFs. Phase 2 runs for 20 timesteps, which was found to be enough time for both activation peaks to fall below the interaction threshold (), but not enough time for the fields to return fully to the resting level. Each simulation runs for a total of 200 timesteps.
@@ 2.2.2 Example simulations: contextual modulation
| adjacency context | low CA (6) low conn (6) have (6) | - | low CA (6) low conn (0.4) have (6) |
| possession context | low CA (6) high conn (6) have (6) | - | low CA (6) low conn (0.4) have (6) |
@@
Examples of simulated interpretation in each context are displayed in Figure 8. In phase 1 (up to timestep 90) the have node becomes active, and begins to send input to the conn field (see Figure 3). The CA field quickly forms a stabilized peak corresponding to an interpretation of low control asymmetry in both conditions. Due to field coupling, this CA peak sends input to the conn field on both the low and high ends (see Figure 4). In the adjacency context, an activation peak corresponding to low conn stabilizes and suppresses the rest of the field via lateral inhibition. In the possession context, the winning activation peak is on the high side of the conn field. During phase 2 (from timesteps 91–110), all activation decreases towards resting level. Activation of the have node returns almost back to its resting level. In the fields, activation falls below the interaction threshold (), but does not return fully to the resting level. At timestep 111 (the beginning of phase 3), field activation is still highly non-uniform, reflecting residual activation from phase 1. In the CA field, another peak forms on the low side of the field in both conditions. In the conn field, the location of the peak differs by condition, despite the fact that the phase 3 inputs are identical between conditions. In the adjacency context, phase 3 (target sentence) processing stabilizes on another interpretation of low conn; in the possession context, the stabilized peak corresponds to an interpretation of high conn. In this way, the evolution of activation in phase 3 is shaped by the preceding state of the system from phase 1.
2.2.3 1000 simulations per condition
@@
The examples in Figure 8 were selected in order to demonstrate the effect of preceding context (phase 1 evolution) on target sentence interpretation (phase 3 evolution). However, not every simulated run of the model exhibits the same effect. The presence of noise in the model introduces a stochastic influence on the location of field stabilization in each simulation. Moreover, the weak low conn input in phase 3 introduces an overall bias for low connectedness in both contexts. In order to examine the robustness of the contextual modulation effect, we simulated 1000 instances of interpretation in each of the two contexts. As seen in Figure 9, there is a bimodal distribution of interpretations in both contexts. That is, for each context, an interpretation corresponding to either adjacency or possession for the target sentence was possible. In this way, the behavior of the model can be qualitatively described as corresponding to discrete outcomes, despite the underlyingly continuous state space (continuous features, continuous time, continuous activation). The apparent discreteness of the interpretations of have is consistent with the intuitions of experiment participants (Zhang, 2021) and linguistic researchers (Myler, 2016). The discreteness arises from the structure imposed by the coupling patterns (node-field and field-field), along with the nonlinear activation dynamics which drive selection of a particular field location on each simulation. Nonetheless, there is gradient variability in the precise location of each activation peak, due to noise in each of the model components.
Importantly, the likelihood of each interpretation (low conn or high conn) was influenced by context. In the adjacency context, low conn (adjacency) interpretations were much more likely, due to the combined influence of the context and the weak bias for low conn coming from the nouns in phase 3. In the possession context, high conn (possession) and low conn (adjacency) interpretations were approximately equally likely, due to the competing influences from the context (biased towards high conn) and from the bias for low conn coming from the nouns. Thus, while context does not completely determine the course of field evolution, it exerts a strong enough influence to be observable over many simulations.
2.3 Simulated acceptability
How do we relate these simulated results to the observed acceptability results described in Section 1? We propose that acceptability judgments are related to the distance between the activation peak in a neural field and the centroid of the distribution of input to the neural field. In particular, we define acceptability as the reciprocal of one plus the distance between the location of the activation peak and the centroid of the input distribution, as in Eq. 7.
Thus, acceptability ranges from to , where the size of the field. Eq. 7 indexes the degree to which an activation peak is consistent with “expected” activation under this set of inputs, operationalized as the centroid of the input distribution. An actual activation peak at a given time can deviate from the “expected” activation peak because of the influence of context and noise. The degree to which the activation peak deviates from the input centroid measures the degree to which an interpretation deviates from a canonical or expected interpretation. We focus here on deviation in the conn field, because the CA field behaves similarly across all simulations under this parameter set. In our simulation of a canonical alienable possession interpretation of have (Figure 7), all of the inputs to the conn field are centered at , the center of the field (mid conn). Thus, the centroid of the input to the conn field is . As mentioned above, under this set of inputs, the model will always tend to form an activation peak near the center of the field, leading to high acceptability. On that particular simulation, the location of the peak was , leading to acceptability .
Now consider our simulations of a sentence with have and two unrelated, inanimate arguments (Figure 8). In this case, the centroid of the input to the conn field during the target sentence is slightly to the left of the center of the field (), due to the weak low conn input coming from the nouns. Thus, while both an adjacency interpretation (peak at low conn) and an inalienable possession interpretation (peak at high conn) will lead to relatively low acceptability (since both peaks will be relatively distant from the input centroid), adjacency interpretations will tend to have higher acceptability than inalienable possession interpretations, because the input centroid is slightly to the left of the center, towards the adjacency end of the field. For example, in the simulations in Figure 8, acceptability in the adjacency context was 0.047, and acceptability in the possession context was 0.041 (cf. 0.25 for the canonical alienable possession interpretation in Figure 7). These observations—overall low acceptability of have sentences with two unrelated, inanimate arguments, but increased acceptability when preceded by an adjacency reading of have relative to an inalienable possession reading of have—are consistent with measured acceptability ratings (Zhang et al., 2018). Figure 10 displays the distribution of simulated acceptability across the same simulations reported in Figure 9.
It can be seen that there is a high degree of overlap between simulated acceptability ratings in each context, because within each of the regions of the conn field (adjacency or possession), random variability can move the activation peak closer to or further from the input centroid. However, the distribution of acceptability is shifted higher overall in the adjacency context relative to the possession context, because activation peaks on the left side of the field tend to be closer to the input centroid than peaks on the right side of the field. To illustrate the by-context difference in acceptability more clearly, Figure 11 plots mean simulated acceptability by context, which is consistent with the empirical observation (Zhang et al., 2018) that mean acceptability is higher in the adjacency context relative to the possession context.
We note that, while the definition of acceptability in Eq. 7 is a heuristic, rather than a genuine neural dynamic account of acceptability judgments, the definition is generalizable. Using Eq. 7, one could calculate acceptability from any set of field inputs (corresponding to the meaning of a sentence) and any neural activation peak (corresponding to a particular interpretation of a sentence). Moreover, Eq. 7 may suggest avenues for implementing a neural dynamic model of acceptability judgments, in that the behavior of such a model should exhibit the relation defined here (i.e., higher acceptability corresponds to lower deviation between an activation peak and the input centroid). Analogous neural mechanisms may be found in the domain of movement error detection. For example, in the DIVA model of speech production (Tourville and Guenther, 2011), predicted sensory consequences of motor commands (“efference copies”: von Holst and Mittelstaedt, 1950) are compared to incoming sensory signals. A mismatch between predicted and actual sensation leads to detection of an error, and a corresponding increase in neural activation (Tourville et al., 2008). Intuitions about sentence acceptability may arise from a similar neural mechanism which compares an expected interpretation to an actual interpretation. Such a mechanism may underlie the electrophysiological difference measured between have processing in the two contexts (Zhang et al., 2018). We leave it to future work to flesh out the details of this mechanism.
2.4 Simulated individual variation